OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5 A. y – 1 = 5(x + 9) B. y – 9 = 5(x + 1) C. y – 9 = 5(x – 1) D. y + 9 = 5(x – 1)

OpenStudy (nurali):

y-y1=m(x-x1)

OpenStudy (nurali):

(x1,y1)=(1,9) m=5

OpenStudy (nurali):

put in formula and tell me what answer?

OpenStudy (nurali):

reply @sapphire13

OpenStudy (nurali):

????

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Well im guessing its C.

OpenStudy (anonymous):

this is too hard???

OpenStudy (anonymous):

can you help @tswavy

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5 A. y + 4 = 5x – 10 B. y + 4 = 5(x + 2) C. y – 4 = 5(x – 2) D. y – 4 = 5(x + 2)

OpenStudy (anonymous):

use the equation y-y1=m(x-x1)

OpenStudy (anonymous):

@sapphire13 the answer is y-9=m(x-1)

OpenStudy (anonymous):

and for your second question the answer is y-4=5(x+2)

OpenStudy (anonymous):

Thanks @tswavy

OpenStudy (anonymous):

You're welcome (: Hope it helped! @sapphire13

OpenStudy (anonymous):

It did :)